# Inverse sampling intensity weighting for preferential sampling adjustment

### v1

@@ Abstract
Traditional geostatistical methods assume independence between observation locations and the spatial process of interest. Violations of this independence assumption are referred to as preferential sampling (PS). Standard methods to address PS rely on estimating complex shared latent variable models and can be difficult to apply in practice. We study the use of inverse sampling intensity weighting (ISIW) for PS adjustment in model-based geostatistics. ISIW is a two-stage approach wherein we estimate the sampling intensity of the observation locations then define intensity-based weights within a weighted likelihood adjustment. Prediction follows by substituting the adjusted parameter estimates within a kriging framework. A primary contribution was to implement ISIW by means of the Vecchia approximation, which provides large computational gains and improvements in predictive accuracy. Interestingly, we found that accurate parameter estimation had little correlation with predictive performance, raising questions about the conditions and parameter choices driving optimal implementation of kriging-based predictors under PS. Our work highlights the potential of ISIW to adjust for PS in an intuitive, fast, and effective manner.
Keywords: model-based geostatistics, preferential sampling, kriging, point process, inverse weighting
@@ 1 Introduction
The field of geostatistics comprises a set of methods to infer properties and predict unknown values (or functions) of a spatially continuous process $\{S(\textbf{x}):\textbf{x}\in\mathcal{D}\subseteq\mathbb{R}^{2}\}$ from a discrete set of observation locations, denoted $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{n}\}$. Model-based approaches estimated via maximum likelihood estimation (MLE) typically assume $S$ is a realization from a random stochastic process, and subsequent inference and prediction treat X as fixed (Diggle et al.,, 1998). This assumption further implies that the sampled locations are independent of the spatial process of interest, or in symbols, $[\textbf{X},S]=[\textbf{X}][S]$, where square brackets indicate the probability distribution. Violations of this independence assumption are referred to as preferential sampling (PS). There is strong evidence to suggest neglecting PS can negatively impact both geostatistical inference and prediction (Diggle et al.,, 2010; Pati et al.,, 2011; Gelfand et al.,, 2012).
Early works in PS examined X under the lens of a marked point process. Schlather et al., (2004) designed two Monte Carlo tests to detect dependence between marks and their locations. Ho and Stoyan, (2008) introduced the classic PS adjustment model, referred to here as the shared latent process (SLP) model, and derived several of its properties. Full estimation and prediction procedures for the SLP model were formally established in the seminal work by Diggle et al., (2010).
@@
$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ [\textbf{X}|S]\sim IPP(\lambda),\\ \lambda(\textbf{x})=\exp(\alpha+\beta S(\textbf{x})),\\ S\sim GP(0,C_{\theta}(\cdot,\cdot)).\\ \end{split}$ (1)
@@
While the SLP framework has been extended to handle covariates in the mean model, non-Gaussian likelihoods, and multiple shared latent processes (Pati et al.,, 2011; Watson et al.,, 2019), for the purposes of illustration, we focus on the original formulation of a constant mean, Gaussian likelihood, and one shared latent process. Under the SLP model, sampling locations are no longer assumed to be fixed and are instead a realization from a log Gaussian Cox process (LGCP). The latent process $S$ is assumed to follow a zero mean second order stationary Gaussian process (GP) and for the remainder of this paper, we further assume $C_{\theta}$ to be an isotropic Matérn covariance function. The SLP takes its name from the presence of $S$ in both the mean model of Y and the intensity function of X. The parameter $\beta$ can be interpreted as a single PS coefficient controlling the preferentiality of the SLP. The sign of $\beta$ determines the direction of the preference, while its magnitude determines the likelihood of sampling extreme values.
@@
Subsequent research on PS methods has remained largely faithful to the SLP model. Pati et al., (2011) extended the SLP model structure to a flexible Bayesian framework estimated by MCMC, and proved the posterior consistency of each of the mean, covariance and PS parameters under increasing domain asymptotics. Gelfand et al., (2012) also used Bayesian estimation, and introduced a framework to compare prediction surfaces under PS between different methods. A surprising finding from their simulation analysis was that inclusion of an informative covariate was not sufficient to correct for predictive bias. This empirical finding highlights the importance of the SLP model even when informative covariates explaining the dependence between X and $S$ are available. Dinsdale and Salibian-Barrera, (2019) significantly improved the computational efficiency of the SLP model relative to the previous simulation-based methods by means of the stochastic partial differential equations (SPDE) approach (Lindgren et al.,, 2011) combined with a Laplace approximation implemented in the Template Model Builder (TMB) R package (Kristensen et al.,, 2016). Watson et al., (2019) also fitted the SLP model using the SPDE approach but with an integrated nested Laplace approximation implemented in the R-INLA software (Rue et al.,, 2009), which enjoys comparable computational gains to the approach in Dinsdale and Salibian-Barrera, (2019). The authors further defined a framework to model PS spatio-temporal data and better emulate the evolution of X over time compared to the original SLP model defined in (1) above.
@@
Several other contributions extend the SLP framework to explore a rich set of important applications, including air pollution monitoring (Lee et al.,, 2011, 2015), species distribution modeling (Manceur and Kühn,, 2014; Fithian et al.,, 2015; Conn et al.,, 2017; Pennino et al.,, 2019; Gelfand and Shirota,, 2019; Fandos et al.,, 2021), disease surveillance (Rinaldi et al.,, 2015; Cecconi et al.,, 2016; Conroy et al.,, 2023), phylodynamic inference (Karcher et al.,, 2016), hedonic modelling (Paci et al.,, 2020), bivariate spatial data (Shirota and Gelfand,, 2022), spatially-varying PS (Amaral et al.,, 2023), optimal design under PS (Ferreira and Gamerman,, 2015), space-filling designs (Ferreira,, 2020; Gray and Evangelou,, 2023), a hypothesis test to detect PS in spatio-temporal data (Watson,, 2021), and exact Bayesian inference for the SLP model (Moreira and Gamerman,, 2022; Moreira et al.,, 2023). The improved predictive ability and parameter estimation within the SLP framework contribute to the popularity of model-based approaches accounting for dependence between spatial processes and their observation locations.
@@
While the benefits of the SLP structure and the ubiquity of PS data support its widespread adoption for spatial data analysis, the SLP model still contains significant limitations. Few software packages and out-of-the-box implementations exist for SLP estimation, often coming in the form of custom INLA or MCMC sampler code. Furthermore, the SLP model is challenging to integrate with modern spatial approximation techniques, many of which have become the standard in spatial data analysis due to their scalability and feasibility, offering computationally efficient alternatives to the MLE while maintaining high accuracy (Heaton et al.,, 2019). Among these methods, only the SPDE approach has a well-documented implementation of a PS solution, and incorporating the SLP model into any other method would take a substantial effort. The SLP model requires joint evaluation of both the response and observation likelihood, not a trivial task. Prediction based on the SLP model also can suffer from high computational cost. While parameter estimation for the SLP model may scale well with the SPDE approach, prediction depends on summarizing the posterior distribution of $[S|\textbf{X},\textbf{Y}]$ for each prediction point, which can be infeasible even on a moderately sized grid for both the Laplace approximation and MCMC.
@@
An alternative strategy to the SLP is to incorporate dependence through the sampling intensities $\lambda$ of the observations, rather than the entire likelihood of X. These methods have been developed to much less fanfare compared to the SLP, and it remains unclear how well they work in geostatistical applications. The two main uses of sampling intensities are as a covariate in the mean model of Y or as inverse weights in a likelihood adjustment (Vedensky et al.,, 2023). A potential advantage of sampling intensity methods is their robustness to the form of PS relative to the SLP framework, which addresses a very specific type of clustering process.
@@
The biggest impediment to the use of the sampling intensity is the need to estimate this intensity from the observed locations, the driving reason such approaches have been sidelined in favor of full model-based approaches. Unlike in survey methodology where survey weights are fixed and known, sampling weights in geostatistical applications are largely unknown to the investigator and must be estimated. Unfortunately, nonparametric kernel smoothing estimators of the intensity (Diggle,, 1985; Berman and Diggle,, 1989) have few theoretical guarantees and are not consistent for the true intensity (Guan,, 2008).
@@
Even so, there is preliminary evidence that methods using nonparametrically estimated sampling intensities can still mitigate the effects of PS and, in some scenarios, can predict a surface better than the SLP approach in practice. Reich and Fuentes, (2012) noted how the Bayesian SLP model parameters could be estimated without MCMC by replacing the shared latent process term in the mean model with some function of the sampling intensity $g(\lambda)$. Since the dependence between X and $S$ is completely contained within $g(\lambda)$, the locations X can once again be treated as fixed, so long as $\lambda$ is estimated well. They show improved prediction using the sampling intensity covariate in a 1-dimensional kriging example. In another example, Zidek et al., (2014) provide an approach using estimated weights for air pollution monitoring. Instead of working in a continuous study region, however, these authors considered a finite superpopulation of possible sampling sites and the probability of selecting any site over time was modeled by a logistic regression. The estimated inverse probabilities were then used as weights in a Horvitz-Thompson style design-based estimator for unbiased estimation of parameters under PS. In a third example, Schliep et al., (2023) also considered a finite superpopulation and similarly deviated from traditional model-based geostatistics by using estimated sampling intensity weights to recover the estimated parameter values had the model been fit on the superpopulation, rather than estimate the parameters of the spatial process $S$. These authors also derived a kriging variance estimate adjusted for PS. Finally, Vedensky et al., (2023) conducted a simulation experiment comparing a univariate marginal composite likelihood (CL) weighted by inverse sampling intensities estimated by the MASS::kde2d function in R to the SLP and unweighted CL, and showed improved performance compared to no adjustment. We will refer to methods using inverse sampling intensities as weights in a weighted likelihood adjustment as inverse sampling intensity weighting (ISIW).
@@
Despite preliminary investigations into ISIW for PS adjustment, specifics regarding the effectiveness of such approaches within model-based geostatistics remains largely unknown. In particular, it is unclear whether they reliably estimate both mean and covariance parameters for a latent spatial process on a 2D continuous surface. Composite likelihoods beyond pairwise difference and univariate marginal have also not been explored for ISIW. Finally, the robustness of ISIW and the SLP to misspecification has not been well-studied. In the sections below, we provide an expanded evaluation of ISIW methods, comparing the performance of MLE, SLP, and ISIW applied to the Vecchia and pairwise marginal CLs across multiple random fields and PS designs.
@@
Our work proceeds as follows. We start off in Section 2 by providing background on model-based geostatistics and introduce the key methods we use as a basis for ISIW. In Section 3, we present the implementation of ISIW in detail. In Section 4, we conduct a comparison analysis of the MLE, SLP, and ISIW through a set of simulation studies. In Section 5 we apply the same methods to the famous Galicia moss dataset which has been the dataset of choice when investigating PS. In Section 6 we finish with final remarks and discuss the implications of our work, and outline future directions for continuing investigation.
@@ 2 Model-based geostatistics
2.1 Estimation
2.1.1 Maximum likelihood estimation
@@
$\begin{split}[\textbf{Y}|\textbf{X},S]\sim N_{n}(\mu\textbf{1}_{n}+S(\textbf{X}),\tau^{2}I_{n}),\\ S\sim GP(0,C_{\boldsymbol{\theta}}(\cdot,\cdot)).\\ \end{split}$
@@
$\begin{split}[\textbf{Y},\textbf{X}]\propto N_{n}(\mu\textbf{1}_{n},\Sigma_{n}(\theta)+\tau^{2}I_{n}),\\ \end{split}$ (2)
@@
Under increasing domain asymptotics, the MLE for $\boldsymbol{\psi}$ is consistent and asymptotically normal (AN) (Mardia and Marshall,, 1984; Bachoc,, 2014, 2020). Results are more challenging under fixed domain asymptotics due to the inclusion of increasingly close observations of a spatially correlated process. Specifically, for a Matérn Gaussian process with known $\nu$, only a subcomponent of the covariance parameters known as the microergodic parameter is consistently estimable and asymptotically normal (Zhang,, 2004; Kaufman and Shaby,, 2013). We refer to this parameter as $\kappa$, given by $\sigma^{2}/\phi^{2\nu}$ for Matérn covariance functions. The nugget $\tau^{2}$ is also estimable (Tang et al.,, 2021) while the variance $\sigma^{2}$ and range $\phi$ are not (Zhang,, 2004). Although the smoothness parameter $\nu$ is theoretically estimable, it is numerically challenging to estimate and is conventionally treated as fixed (Loh et al.,, 2021). Fixed effect coefficients in the mean are typically non-estimable (Wang et al.,, 2020), except under specific smoothness conditions on the covariates (Yu,, 2022; Bolin and Wallin,, 2024; Gilbert et al.,, 2024).
@@
The main drawback for practical use of the MLE is its computational burden. Evaluating the likelihood requires an inversion of the covariance matrix, which has $O(n^{3})$ time and $O(n^{2})$ space complexity making it infeasible for applications for moderately sized $n$. Recent advancement in modern spatial statistics has focused on GP parameter estimation using approximations which massively reduce computational burden in exchange for marginal decreases in efficiency. We explore composite likelihood and the Vecchia approximation, two approaches which have been widely adopted in geostatistics and are particularly conducive to weighting.
2.1.2 Composite likelihood
Sticking with our previous notation, let $\textbf{Y}=(Y_{1},Y_{2},...,Y_{n})^{\top}\in\mathbb{R}^{n}$ be a $n\times 1$ random vector with probability density $f(\textbf{y};\boldsymbol{\psi})$ for unknown $r$-dimensional parameter vector $\boldsymbol{\psi}\in\Psi\subseteq\mathbb{R}^{r}$. Define $\{\mathcal{A}_{1},...,\mathcal{A}_{K}\}$ to be a set of $K$ marginal or conditional events with associated likelihood $\mathcal{L}_{k}(\theta;\textbf{y})\propto f(\textbf{y}\in\mathcal{A}_{k};\boldsymbol{\psi})$. The log composite likelihood (CL) is the weighted sum of the event-specific log likelihoods
where $\textbf{w}:=(w_{1},...,w_{K})$ is a vector of weights, not necessarily non-negative and the density $f$ follows that defined in (2). The maximum CL estimate is defined as $\hat{\boldsymbol{\psi}}_{C}:=\text{argmax}_{\boldsymbol{\psi}}\log\mathcal{L}_{C}(\boldsymbol{\psi};\textbf{y})$. One common choice of CL is the univariate marginal where the log likelihood is the sum of the log marginal density of each observation.
Typically, the use of this univariate marginal in geostatistics is limited because it ignores dependencies between observations and thus cannot estimate covariance parameters. For this reason, pairwise CLs have been much more popular (Varin et al.,, 2011). Bevilacqua and Gaetan, (2015) compared the efficiency of three different pairwise CL estimators: pairwise marginal (PMLE), pairwise conditional (PCMLE), and pairwise difference (PDMLE) CLs. These authors concluded that the PMLE outperformed all other pairwise CLs and recommended its use over the PCMLE and PDMLE. The likelihoods for each pairwise CL are
These authors also proved, under increasing domain asymptotics, the consistency and asymptotic normality of $\boldsymbol{\hat{\psi}}_{C}$ estimated from (3). Bachoc et al., (2019) proved the same properties for the PMLE and PCMLE in the one-dimensional setting under fixed domain asymptotics. However, the true utility of CL lies in its computational speed and robustness to misspecification. While the MLE requires a $O(n^{3})$ matrix inversion and specification of the joint density, pairwise CL consists only of $O(n^{2})$ terms and only requires correct specification of second order densities. The computational efficiency of pairwise CL can be further enhanced by using the weights w to only include pairwise observations within a certain distance $d$ apart (Bevilacqua and Gaetan,, 2015).
In later sections, we will use the PMLE as a basis for ISIW due to its superior performance over other pairwise likelihoods. In particular, we do not consider the univariate marginal or PDMLE because they do not directly estimate the entirety of $\boldsymbol{\psi}$. In the next section, we discuss the Vecchia approximation, an alternative likelihood approach that achieves even greater computational and statistical efficiency than the CL.
2.1.3 Vecchia approximation

### v2

@@ Abstract
Traditional geostatistical methods assume independence between observation locations and the spatial process of interest. Violations of this independence assumption are referred to as preferential sampling (PS). Standard methods to address PS rely on estimating complex shared latent variable models and can be difficult to apply in practice. We study the use of inverse sampling intensity weighting (ISIW) for PS adjustment in model-based geostatistics. ISIW is a two-stage approach wherein we estimate the sampling intensity of the observation locations then define intensity-based weights within a weighted likelihood adjustment. Prediction follows by substituting the adjusted parameter estimates within kriging. We introduce an implementation of ISIW based on the Vecchia approximation, enabling computational gains while maintaining strong predictive accuracy. Interestingly, we found that ISIW outpredicts standard PS methods under misspecification of the sampling design, and that accurate parameter estimation had little correlation with predictive performance, raising questions about the conditions driving optimal implementation of kriging-based predictors under PS. Our work highlights the potential of ISIW to adjust for PS in an intuitive, fast, and effective manner. We illustrate these ideas on spatial prediction of lead concentrations measured through moss biomonitoring data in Galicia, Spain, and $\text{PM}_{2.5}$ concentrations from the U.S. EPA Air Quality System network in California.
Keywords: inverse sampling intensity weighting, ISIW, model-based geostatistics, preferential sampling, kriging, point process
@@ 1 Introduction
The field of geostatistics comprises a set of methods to infer properties and predict unknown values (or functions) of a spatially continuous process $\{S(\textbf{x}):\textbf{x}\in\mathcal{D}\subseteq\mathbb{R}^{2}\}$ from a discrete set of observation locations, denoted $\textbf{X}=\{\textbf{x}_{1},\textbf{x}_{2},\ldots,\textbf{x}_{n}\}$. Model-based approaches estimated via maximum likelihood estimation (MLE) typically assume $S$ is a realization from a random stochastic process, and subsequent inference and prediction treat X as fixed (diggle_model-based_1998). This assumption further implies that the sampled locations are independent of the spatial process of interest, or in symbols, $\textbf{X},S]=[\textbf{X}][S]$, where square brackets indicate the probability distribution. Violations of this independence assumption are referred to as preferential sampling (PS). Several studies have shown that neglecting PS can introduce substantial bias in geostatistical inference and prediction (diggle_geostatistical_2010; pati_bayesian_2011; gelfand_effect_2012).
Early works in PS examined X under the lens of a marked point process. schlather_detecting_2004 designed two Monte Carlo tests to detect dependence between marks and their locations. ho_modelling_2008 introduced the classic PS adjustment model, referred to here as the shared latent process (SLP) model, and derived several of its properties. Full estimation and prediction procedures for the SLP model were formally established in the seminal work by diggle_geostatistical_2010.
@@
[…]
@@
While the SLP framework has been extended to handle covariates in the mean model, non-Gaussian likelihoods, and multiple shared latent processes (pati_bayesian_2011; watson_general_2019), for the purposes of illustration, we focus on the original formulation of a constant mean, Gaussian likelihood, and one shared latent process. Under the SLP model, sampling locations are no longer assumed to be fixed and are instead a realization from a log Gaussian Cox process (LGCP). The latent process $S$ is assumed to follow a zero mean second order stationary Gaussian process (GP). The SLP takes its name from the presence of $S$ in both the mean model of Y and the intensity function of X. The parameter $\beta$ can be interpreted as a single PS coefficient controlling the preferentiality of the SLP. The sign of $\beta$ determines the direction of the preference, while its magnitude determines the likelihood of sampling extreme values.
@@
Subsequent research on PS methods has remained largely faithful to the SLP model. pati_bayesian_2011 extended the SLP model structure to a flexible Bayesian framework estimated by MCMC, and proved the posterior consistency of each of the mean, covariance and PS parameters under increasing domain asymptotics. gelfand_effect_2012 also used Bayesian estimation, and introduced a framework to compare prediction surfaces under PS between different methods. A surprising finding from their simulation analysis was that inclusion of an informative covariate was not sufficient to correct for predictive bias. This empirical finding highlights the importance of the SLP model even when informative covariates explaining the dependence between X and $S$ are available. dinsdale_methods_2019 significantly improved the computational efficiency of the SLP model relative to the previous simulation-based methods by means of the stochastic partial differential equations (SPDE) approach (lindgren_explicit_2011) combined with a Laplace approximation implemented in the Template Model Builder (TMB) R package (kristensen_tmb_2016). watson_general_2019 also fitted the SLP model using the SPDE approach but with an integrated nested Laplace approximation implemented in the R-INLA software (rue_approximate_2009), which enjoys comparable computational gains to the approach in dinsdale_methods_2019. The authors further defined a framework to model PS spatio-temporal data and better emulate the evolution of X over time compared to the original SLP model defined in (1) above.
@@
Several other contributions extend the SLP framework to explore a rich set of important applications, including air pollution monitoring (lee_constructing_2011; lee_impact_2015), species distribution modeling (manceur_inferring_2014; fithian_bias_2015; conn_confronting_2017; pennino_accounting_2019; gelfand_preferential_2019; fandos_dynamic_2021), disease surveillance (rinaldi_sheep_2015; cecconi_preferential_2016; conroy_shared_2023), phylodynamic inference (karcher_quantifying_2016), hedonic modelling (paci_spatial_2020), bivariate spatial data (shirota_preferential_2022), spatially-varying PS (amaral_model-based_2023), optimal design under PS (ferreira_optimal_2015), space-filling designs (ferreira_geostatistics_2020; gray_design_2023), a hypothesis test to detect PS in spatio-temporal data (watson_perceptron_2021), and exact Bayesian inference for the SLP model (moreira_analysis_2022; moreira_presence-only_2024). The SLP’s flexibility in addressing sampling bias has driven the popularity of model-based approaches that account for the inherent dependence between spatial processes and their observation locations.
@@
Despite these advantages, the SLP framework contains significant limitations that complicate its application. Few software packages and out-of-the-box implementations exist for SLP estimation, often coming in the form of custom INLA or MCMC sampler code. Furthermore, the SLP model is challenging to integrate with modern spatial Gaussian process approximation techniques, many of which have become the standard in spatial data analysis due to their scalability and feasibility, offering computationally efficient alternatives to the MLE while maintaining high accuracy (heaton_case_2019). Among these methods, only the SPDE approach has a well-documented implementation of a PS solution, and incorporating the SLP model into any other method would take a substantial effort. The SLP model requires the nontrivial task of jointly modeling both the response and observation likelihood. Prediction based on the SLP model can also suffer from high computational cost. While parameter estimation for the SLP model may scale well with the SPDE approach, prediction depends on summarizing the posterior distribution of $S|\textbf{X},\textbf{Y}]$ for each prediction point, which can be infeasible even on a moderately sized grid for both the Laplace approximation and MCMC.
@@
An alternative strategy to the SLP is to incorporate dependence through the sampling intensities $\lambda$ of the observations, rather than the entire likelihood of X. The two main potential uses of sampling intensities are as a covariate in the mean model of Y or as inverse weights in a likelihood adjustment (vedensky_look_2023). It remains unclear how well these methods work in geostatistical applications. A potential advantage of sampling intensity methods is their robustness to the form of PS relative to the SLP framework, which requires full parametric specification of the likelihood of $\mathbf{X}$.
@@
The biggest impediment to the use of the sampling intensity is the need to estimate this intensity from the observed locations. Unlike in survey methodology where survey weights are fixed and known, sampling weights in geostatistical applications are largely unknown to the investigator and must be estimated. Unfortunately, nonparametric kernel smoothing estimators of the intensity (diggle_kernel_1985; berman_estimating_1989) have few theoretical guarantees and are not consistent for the true intensity without knowledge of spatial covariates (guan_consistent_2008).
@@
Even so, there is preliminary evidence that methods using nonparametrically estimated sampling intensities can still mitigate the effects of PS. mateu_accounting_2012 noted how the Bayesian SLP model parameters could be estimated without MCMC by replacing the shared latent process term in the mean model with some known function of the sampling intensity $g(\lambda)$. In another example, zidek_reducing_2014 provide an approach using estimated weights for air pollution monitoring. Instead of working in a continuous study region, however, these authors considered a finite superpopulation of possible sampling sites and the probability of selecting any site over time was modeled by a logistic regression. The estimated inverse probabilities were then used as weights in a Horvitz-Thompson style design-based estimator for unbiased estimation of parameters under PS. schliep_correcting_2023 also considered a finite superpopulation and similarly deviated from traditional model-based geostatistics by using estimated sampling intensity weights to recover the estimated parameter values and kriging variance had the model been fit on the superpopulation, rather than estimate the parameters of the spatial process $S$. Finally, vedensky_look_2023 conducted a simulation experiment comparing a univariate marginal composite likelihood (CL) weighted by inverse sampling intensities estimated by the MASS::kde2d function in R to the SLP and unweighted CL, and showed improved performance compared to no adjustment. We will refer to methods using inverse sampling intensities as weights in a weighted likelihood adjustment as inverse sampling intensity weighting (ISIW).
@@
Despite preliminary investigations into ISIW for PS adjustment, specifics regarding the effectiveness of such approaches within model-based geostatistics remain largely unknown. In particular, it is unclear whether they reliably estimate both mean and covariance parameters for a latent spatial process on a 2D continuous surface. Composite likelihoods beyond pairwise difference and univariate marginal have also not been explored for ISIW. Finally, the robustness of ISIW and the SLP to misspecification has not been well-studied. In the sections below, we provide an expanded evaluation of ISIW methods, comparing the performance of MLE, SLP, and ISIW across multiple random fields and PS designs.
@@
Our work proceeds as follows. We start off in Section 2 by providing background on model-based geostatistics and introduce the key methods we use as a basis for ISIW. In Section 3, we present the implementation of ISIW in detail. In Section 4, we conduct a comparison analysis of the MLE, SLP, and ISIW through a set of simulation studies. In Section 5 we apply the same methods to the famous Galicia moss dataset which has been the dataset of choice when investigating PS, as well as a dataset of air pollution measurements taken from the Air Quality System (AQS) in California. In Section 6 we finish with final remarks and discuss the implications of our work, and outline future directions for continuing investigation.
@@ 2 Model-based geostatistics
2.1 Maximum likelihood estimation
@@
[…]
@@
[…]
@@
The main drawback for practical use of the MLE is its computational burden. Evaluating the likelihood requires an inversion of the covariance matrix, which has $O(n^{3})$ time and $O(n^{2})$ space complexity making it infeasible for applications for moderately sized $n$. Recent advancement in modern spatial statistics has focused on GP parameter estimation using approximations which massively reduce computational burden in exchange for marginal decreases in efficiency. In the next section, we discuss the Vecchia approximation, one such alternative likelihood approach that achieves high computational and statistical efficiency.
2.2 Vecchia approximation